# KIT’s Low-resource Speech Translation Systems for IWSLT2025: System Enhancement with Synthetic Data and Model Regularization

### v1

@@ 1 Introduction
In this paper, we present our submissions to the IWSLT 2025 low-resource track. We participate in three language pairs, translating from Bemba (ISO: bem), North Levantine Arabic (ISO: apc), and Tunisian Arabic (ISO: aeb) into English. Our approach follows the unconstrained track, reflecting practical scenarios by leveraging all available resources, including multilingual pre-trained models and external datasets.
@@ 3.2 Model Regularization
Regularization remains a simple yet powerful way to boost the generalisation capacity of neural sequence models, and has already proved valuable in machine translation through techniques such as R-Drop and its variants Wu et al. (2021); Xu et al. (2022). Motivated by the recent success of intra-distillation (ID) in low-resource MT Romney Robinson et al. (2024), we extend ID to all three tasks: ASR, MT, and ST, based on the public implementation with the following modification [^1] https://github.com/fe1ixxu/Intra-Distillation/.
Unlike previous work that directly fine-tunes a pre-trained model with a loss that combines the task objective and ID, we notice that direct fine-tuning leads to suboptimal performance in preliminary experiments. We therefore adopt a two-stage approach: (1) vanilla fine-tuning to adapt the pre-trained model to the downstream task, followed by (2) ID fine-tuning to regularize the adapted model with its own intermediate predictions. This simple approach retains the advantages of task-specific adaptation while unlocking the additional robustness that ID provides.

### v2

@@ 1 Introduction
In this paper, we present our submissions to the IWSLT 2025 low-resource track [^1] https://github.com/ZL-KA/IWSLT25-low-resource-KIT. We participate in three language pairs, translating from Bemba (ISO: bem), North Levantine Arabic (ISO: apc), and Tunisian Arabic (ISO: aeb) into English. Our approach follows the unconstrained track, reflecting practical scenarios by leveraging all available resources, including multilingual pre-trained models and external datasets.
@@ 3.2 Model Regularization
Regularization remains a simple yet powerful way to boost the generalisation capacity of neural sequence models, and has already proved valuable in machine translation through techniques such as R‑Drop and its variants Wu et al. (2021); Xu et al. (2022). Motivated by the recent success of intra‑distillation (ID) in low-resource MT Romney Robinson et al. (2024), we extend ID to all three tasks: ASR, MT, and ST, based on the public implementation with the following modification [^2] https://github.com/fe1ixxu/Intra-Distillation/.
Unlike previous work that directly fine-tunes a pre‑trained model with a loss that combines the task objective and ID, we notice that direct fine-tuning leads to suboptimal performance in preliminary experiments. We therefore adopt a two‑stage approach: (1) vanilla fine‑tuning to adapt the pre‑trained model to the downstream task, followed by (2) ID fine‑tuning to regularize the adapted model with its own intermediate predictions. This simple approach retains the advantages of task‑specific adaptation while unlocking the additional robustness that ID provides.